# High Frequency Quoting Under Liquidity Constraints

### v1

@@ 13 Discussion
This section synthesizes the empirical findings presented in the preceding sections, with particular emphasis on the comparative behavior of reference leg selection rules under realistic quoting and execution scenarios. The evaluation framework was structured to isolate and test three competing decision mechanisms: the timestamp-based forecast rule (Hawkes-driven), the state-based CLF rule, and the baseline benchmark rule.
The forecasting rule derived from Hawkes-based arrival modeling demonstrated clear statistical advantage in periods with stable excitation geometry. This was most evident in the reduced slippage profile during episodes where trade initiations displayed directional clustering. The model’s ability to forecast the relative aggressiveness of future events allowed it to anticipate where liquidity would first disappear, and hence quote accordingly. However, this advantage was attenuated in periods where the order flow exhibited higher randomness or event-type inversion, reducing the predictive signal embedded in the arrival ratio.
In contrast, the CLF rule—by construction—remained robust to such statistical variations. Its performance was governed primarily by the instantaneous imbalance in executable quantity across price levels, making it a locally optimal, myopic strategy. While it did not benefit from future information encoded in event history, it incurred less variance across market regimes and was notably more stable under low-activity conditions. The deterministic nature of the CLF rule made it computationally tractable and deployable without a learning phase or model calibration.
Across all experiments, the benchmark strategy served as a neutral baseline, confirming that naive quoting can lead to systematically higher slippage, particularly when spread compression or one-sided flow is present.
Modeling Remark.  The CLF rule is fully deterministic and real-time computable, requiring only an order book snapshot at decision time.
Importantly, our results point to a broader distinction in modeling philosophy: the Hawkes-based rule exploits temporal structure and historical excitation, whereas the CLF rule is anchored in instantaneous state asymmetry. This distinction reflects an implicit divide between modeling order flow memory (i.e., endogenous clustering) and modeling latent liquidity (i.e., revealed imbalance). The differing slippage behavior of these rules, therefore, encodes more than just execution performance—it also offers insight into how informational content is embedded in different layers of market microstructure.
That said, each model has limitations. The Hawkes-based forecast rule requires reliable estimation of excitation parameters, and may underperform in thin markets where event sparsity limits inference. The CLF rule, while robust, may miss cross-contract signals that are only apparent in temporally coupled dynamics.
We now consolidate these results to outline actionable implications for execution strategy design.
@@ 14 Conclusion
This paper examined the impact of reference leg selection on execution quality in calendar spread quoting strategies. By evaluating two distinct decision frameworks—one based on historical excitation patterns (via Hawkes arrival ratios) and the other on contemporaneous order book state (via the Composite Liquidity Factor)—we demonstrated that reference leg choice can materially affect slippage and fill quality, even when quoting spreads are symmetric.
The Hawkes-based forecasting rule proved effective in markets exhibiting persistent excitation geometry, capturing directional order flow and anticipating the side where liquidity would thin out first. In contrast, the CLF rule offered robustness and immediate computability, with stable performance across market regimes, particularly in sparse or noisy environments.
The experimental results reveal that no single quoting strategy dominates across all conditions. Rather, execution performance depends on the interaction between order flow memory and real-time liquidity imbalance. This underscores the need for adaptive quoting logic that blends historical event structure with instantaneous state features.
Future work can explore hybrid strategies that incorporate both Hawkes-based inference and CLF signals, possibly using reinforcement learning to dynamically arbitrate between them. Additionally, incorporating additional market context—such as contract-specific volatility or macroeconomic event markers—may improve the decision framework’s resilience.

### v2

@@ 6 Discussion
The prevailing literature on market microstructure supports such state-dependent performance. Empirical studies show that LOB shape, depth profiles and queue imbalance near the best quotes carry information about short horizon price impact [10, 18]. The CLF is computed directly from limit order book primitives: it responds to instantaneous depth, queue imbalance, and local LOB shape. This design allows it to adjust quickly when the relative liquidity of the two legs changes over short horizons. At the same time, during periods of intense trading activity, high-frequency quotes and trades are subject to microstructure noise and fleeting orders. Hasbrouck [27] emphasises that low-latency trading can amplify such transitory effects in the observed LOB. Under such conditions, a purely cross-sectional LOB-based statistic may produce spurious or short-lived signals.
The Hawkes arrival-ratio, in contrast, is a low-dimensional summary of the event-time dynamics that influence price-impacting events. It aggregates information about how often price-impacting events arrive on each leg, with an explicit memory structure. Under standard Hawkes specifications, intensity forecasts are driven by self- and cross-excitation in recent order-flow [4]. This construction makes the arrival ratio relatively insensitive to very short-lived fluctuations in the order-flow, and better suited to detecting and exploiting persistent directional regimes in the benchmark.
The pattern in benchmark persistence aligns with this microstructure interpretation. In segments with low benchmark persistence, where the oracle switches legs frequently or after only short bursts, CLF4 achieves higher agreement with the oracle than the Hawkes arrival ratio. In the highest-persistence bin, where the benchmark exhibits long, stable regimes, the Hawkes arrival ratio attains the higher agreement score. Collapsing the bins into low-persistence (runs of length at most 58 windows) and high-persistence (runs of at least 59 windows) segments reproduces the aggregate pattern discussed above: in the low-persistence segment CLF4 attains an agreement score of $0.5155$ compared with $0.3903$ for the Hawkes arrival ratio, whereas in the high-persistence segment the Hawkes arrival ratio reaches $0.7060$ compared with $0.6441$ for CLF4. In this sense, the two rules exploit complementary parts of the benchmark’s temporal structure: short-memory cross-sectional signals versus longer-memory event-history signals.
@@ 7 Conclusion
This study develops and evaluates a diagnostic framework for reference-leg selection in multi-contract quoting. Using tick-by-tick NIFTY index futures data, we construct a market-optimal oracle based on realised slippage and compare two reference-leg decision rules: a Hawkes arrival-ratio rule built from reference-impacting order-flow, and a CLF-based rule constructed from contemporaneous LOB shape. In our illustrative application, the Hawkes arrival ratio attains higher agreement with the oracle in high-persistence benchmark regimes, while the CLF-based rule is relatively more accurate when the benchmark switches frequently; in addition, there is a non-trivial set of windows in which both rules match the oracle, indicating that event-history and LOB-state information often point to the same market-optimal benchmark.
Methodologically, the framework is diagnostic in scope. It is parsimonious at runtime and portable across tick-driven, multi-contract markets where quoting is employed,since it relies only on tick-time event histories and LOB snapshots.The diagnostic emphasis is deliberate: it isolates decision quality from strategy-specific implementation details and clarifies where event-history and LOB-state information each contribute to reference-leg selection. Evaluation is anchored by an infeasible hindsight lower bound on deviation from the target spread, which provides a common yardstick for comparing heterogeneous rules, and all tests follow a walk-forward, out-of-sample protocol that relies on model-comparison procedures controlling for data-snooping risk [25, 42].
The construction of the CLF places the framework within the strand of microstructure research that links LOB state to short-horizon execution risk. The incorporation of cross-level log-quote slopes is motivated by Hasbrouck’s analysis of low-latency order books, where steep slopes in the best quotes are associated with fragile depth and heightened sensitivity to incoming orders [27]. In this paper these ingredients are aggregated into a single, direction-sensitive score per leg, designed to be computed from the LOB without repeated high-dimensional re-estimation.
The Hawkes arrival ratio connects the study to intensity-based models of execution risk with limit orders. The prevailing literature investigates limit-order fills and price moves modeled by point-process intensities [23, 14, 24]. The associated execution and price risk is managed by steering these intensities rather than by reacting solely to static book configurations. However, these models are constructed for single contract order-flow and do not naturally extend to multi-contract quoting frameworks and the dependent nature of execution risk in such a setting. The arrival ratio used here is a reduced-form summary of such intensity forecasts, the framework uses their relative magnitudes to choose the leg that is less exposed to future price-impacting flow. In our illustrative application, the non-trivial intersection of windows in which both CLF and Hawkes arrival ratios match the oracle suggests that cross-sectional book shape and short-run order-flow clustering often point to the same market-optimal oracle.
Looking ahead, the same diagnostic structure can be embedded in an explicit control formulation. In such a setting, each quoting rule induces a mapping from an event-time state (summarising recent order-flow, LOB shapes or other market co-variates) to a reference-contract choice. The oracle benchmark provides a natural loss functional based on slippage relative to the choice of market optimal reference contract. Formalising this extension is beyond the scope of the present methodological study, but the results here suggest that combining event-history and LOB-state signals in such a control framework is a promising direction for future work.